OpenStudy (anonymous):

is 5 over six equal to8 over 10?

OpenStudy (saifoo.khan):

\[\frac46=\frac{8}{10}\]Cross multiply.

OpenStudy (saifoo.khan):

\[4 \times 10 = 8 \times 6\]

myininaya (myininaya):

\[\frac{5}{6} \neq \frac{8}{10}; \frac{5}{6} \approx .83; \frac{8}{10}=.8;\]

myininaya (myininaya):

\[\frac{5}{6}>\frac{8}{10}\]